OpenStudy (anonymous):

The formula K(C)= C+ 273 converts Celsius temperature to Kelvin. The formula C(F)= 5/9(F- 32) converts Fahrenheit temperature to Celsius. a) Write a composite function that will convert Fahrenheit temperature to Kelvin. b) Convert the boiling point of water (212 degrees Fahrenheit) and the freezing point of water (32 degrees Fahrenheit) to Kelvin.

OpenStudy (anonymous):

a) To convert Fahrenheit to Kelvin, we can convert it to Celsius first, and then Kelvin. To do this, we can fill in the Celsius function as the missing "variable" of the Kelvin function. K(C(F)) = C + 273 K(C(F)) = C(F) + 273 * K(C(F)) = (5/9)(F-32) + 273 * That would be the composite function that converts Fahrenheit to Kelvin. Also, the Kelvin function should end with 273.15, but I guess it doesn't matter since it's so close. b) For this, we just have to fill in the Fahrenheit number in to the composite function we got in the previous answer. I'll let you do this since it's simple. If you still need help, just ask.

OpenStudy (anonymous):

thank you @ColossalMagnitude

OpenStudy (anonymous):

You welcome. Did you figure out the answer to b), or do you need help with that?

OpenStudy (anonymous):

the first one is 373 kelvin and the second if 273 kelvin. Correct? @ColossalMagnitude

OpenStudy (anonymous):

YUP! Great job!